A water vapour thermometer is calibrated at the ice and steam points and the saturation
vapour pressures (s.v.p) are measured. The saturation vapour pressure Pθ is given by
Pθ =P0(1+αθ+βθ^2) , where P0 is the s.v.p at the ice point, θ is the temperature in ˚C on
the Celsius scale, α and β are constants and α = 40 β ˚C. Determine the temperature of a
medium on the water vapour scale if its temperature on the Celsius scale is 50 ˚C.
1
Expert's answer
2020-07-06T15:25:47-0400
"T=\frac{p_{\theta}}{p_{triple \:point}}T_{triple\: point}"

Thus,

"T=\frac{p_o(1+40\beta\theta+\beta\theta^2)}{0.00603659}0.01^\circ C"

Where, "\theta=50^\circ C" ,Hence

"T=\frac{10p_o(1+4500\beta)}{603659} \: ^\circ C"